For a specific volume of 0.2 m3/kg, find the quality of steam if the absolute pressure is (a) 40 kPa and ( b ) 630 kPa. What is the temperature of each case?
1
Expert's answer
2017-03-15T13:04:05-0400
a) We use the table of Saturated water— Pressure, we calculate the quality as follows:
v = vf + x (vg − vf),
0.2 = 0.001 + x (3.993 − 0.001)
x=0.04985

The temperature is found in Table: T = 75.9°C.
b) Using the values at 0.6 and 0.8 MPa we have:
vg = (0.03 0.2) ×(0.2404 − 0.3157) + 0.3157 = 0.3044
vf=0.0011
0.2 = 0.0011 + x (0.3044 − 0.0011)
x=0.6558
T = (0.03 0.2) × (170.4 − 158.9) + 158.9 = 160.6°C